OpenStudy (anonymous):

2x-3=2(5+x) is this conditional, contradiction or identity? then describe

OpenStudy (saifoo.khan):

dnt knw!

OpenStudy (anonymous):

Start by trying to get the two sides of the equal sign to be of the same form.

OpenStudy (anonymous):

i.e. distribute the 2 to each of the terms in (5+x)

OpenStudy (anonymous):

I have a hard time solving these as it is Im learning though

OpenStudy (anonymous):

Stop that

OpenStudy (anonymous):

contradiction

OpenStudy (anonymous):

I think it is a contradiction.

OpenStudy (anonymous):

Yeah, it's a contradiction because -3 doesn't equal 10

OpenStudy (anonymous):

I think you all should be a little quieter and actually help her solve the problem.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

polpak, it's okay, we're helping her.

OpenStudy (anonymous):

Roseanne. Do you understand any of what they said?

OpenStudy (anonymous):

so polpak can u help me a lil and show me how to solve these I need to learn

OpenStudy (anonymous):

If she has any questions, she can always ask...

OpenStudy (anonymous):

It was 2x - 3, not 2x+3

OpenStudy (anonymous):

She did ask. And all she got was a pointless one word response that was WRONG the first two times. Perhaps you could start from the position of helping rather than just trying to get the answer out quickly and she might not have to ask the next time.

OpenStudy (anonymous):

Hey, I came in late...sorry?

OpenStudy (anonymous):

2x - 3 = 10 + 2x [I distributed the 2 on the left side.] 2x = 13 + 2x [I added 3 to both sides.] 0 = 13 [I subtracted 2x from both sides.] But 0 does not equal 13, this makes it a contradiction.

OpenStudy (anonymous):

so there asking to describe the soulution la is what u just mean describing it?

OpenStudy (anonymous):

LA-MrC, I question your technique

OpenStudy (anonymous):

It's really a one step problem....Just subtract 2x from both sides

OpenStudy (anonymous):

ok im confused

OpenStudy (anonymous):

Then you have to know what conditional, contradiction, or Identity means

OpenStudy (anonymous):

i know what identity means

OpenStudy (anonymous):

Roseanne, I have posted a new question for you.

OpenStudy (anonymous):

Conditional means only one solution, Contradiction means no solutions, Identity means infinite solutions

OpenStudy (anonymous):

condratiction i think means makes no sense like 104 = 5 right? no varaible or am Iwrong

OpenStudy (anonymous):

Yes, correct

OpenStudy (anonymous):

Conditional doesn't mean only one solution.

OpenStudy (anonymous):

It means limited solutions

OpenStudy (anonymous):

Not even that.

OpenStudy (anonymous):

what does it mean

OpenStudy (anonymous):

Contradictory means it is never true (independent of the variables) Conditional means that is is sometimes true (depends on the values of the variables) Identity means that it is always true (indepenent of the variables)

OpenStudy (anonymous):

ok